Suppose we would like to test H_0: μ x = μ y .

Denote the unpaired two-sample normalized test statistic as t 1 , and the ANOVA F test statistic as \\ t 2 .

Suppose for given sample data of equal sample size, we get t 1 = 10 . Compute t 2 .

Hint: For equal sample size, establish a relationship between unpaired two-sample test and ANOVA F test.
Answer #1

Sol:

For unpaired t test (2-sample indpendent t test),

(t)^2=F

here F=t2=t1^2=10^2=100

ANSWER:

t 2 =100